# Still More Shades of Null: A Benchmark for Responsible Missing Value Imputation

### v1

@@ 4.4. Imputation Quality and Fairness
To better understand the impact of different MVI techniques on model performance, we measure the imputation quality of each imputer before training ML models on the imputed training set. The quality of imputation is evaluated across three dimensions: correctness (using F1 for categorical features and RMSE for numerical features), distributional alignment (using KL divergence of imputed values), and fairness (measuring differences in F1, RMSE, and KL divergence between disadvantaged and privileged groups). The higher values of F1 are the better, the lower values of RMSE and KL divergence are the better, and values of imputation fairness closer to zero are the better.
To streamline exposition, we selected one categorical and one numerical feature for each dataset based on their Spearman correlation with the target. However, our analysis takes into account the imputation quality measured on all features with nulls. Importantly, we do not assess imputation quality for boostclean, since this is a joint cleaning and model training method that does not provide imputed values, which we could use to measure imputation quality.
4.4.1. Correctness of imputation
Figures 18 and 19 in Appendix C.1 demonstrate the correctness of imputation of all MVI techniques, except boostclean, for all datasets. From Figure 18, we can notice that one of the ML and DL-based imputers (miss-forest, auto-ml, datawig) is the best in terms of F1 for categorical columns under MCAR and MAR missingness mechanisms. Additionally, Figure 19 shows that all the three ML and DL-based imputers are dominant in terms of RMSE for numerical columns compared to statistical MVI techniques. However, these significant differences in F1 and RMSE do not translate into substantial differences in downstream model performance, as measured by F1 in Section 4.1. This observation further supports our statement made in Section 5.4 that “imputation quality metrics are not very meaningful because a bad data imputation method can be overcome by a strong learner, and so bad performance on imputation quality does not necessarily mean than the MVI will lead to poor predictive performance.”
4.4.2. Distributional alignment of imputation
Figures 20 and 22 in Appendix C.1 illustrate the distributional alignment of imputation of all MVI techniques except boostclean, for all datasets. We find that ML and DL-based imputers, except clustering, outperform statistical approaches in terms of distributional alignment across all missingness mechanisms. The plots clearly demonstrate that, as we highlight in Section 5.4, distributional metrics often serve as more reliable indicators than discrepancy metrics. This is because the distributional alignment of imputed features can significantly influence model training outcomes, emphasizing that assessing only the correctness of imputation is often insufficient.
4.4.3. Fairness of imputation
Figures 21, 23, 24, and 25 in Appendix C.1 show the fairness of imputation of all MVI techniques, except boostclean, for all 6 datasets. From Figures 21 and 23, we observe that statistical MVI techniques are prone to worsen imputation fairness rather than improve it under all missingness mechanisms. Further, comparing imputation fairness with the fairness of the downstream predictive model, discussed in Section 4.2, we can state that good imputation fairness may not lead to model fairness downstream. For example, for the folk dataset, miss-forest was better than median-mode in terms of imputation fairness (F1 score difference of 0.08), yet the downstream model trained on miss-forest-imputed data was less fair than the model trained on median-mode-imputed data (0.03 TRPD).
5. Multi-mechanism Missingness
We now evaluate different MVI techniques under multi-mechanism missingness (Zhou et al., 2024), when MCAR, MAR and MNAR co-exist in both train and test data (S10 in Table 1). This is a more realistic setting that is likely to be encountered in practice (Graham, 2009). We inject 10% of nulls from each mechanism into both train and test sets, for a total of 30% missingness, as in S1-3. In Section 4, we reported results for the best performing ML models for each dataset, which was the Random Forest for diabetes. In this experiment, we focus on diabetes and dive deeper into the performance of different model types.
(a) F1
(b) True Positive Rate Difference (Fairness)
(c) Label stability
Figure 6. Performance of MVI techniques under multi-mechanism missingness on the diabetes dataset, for different model types: Decision Tree (dt__\__clf), Logistic Regression (lr__\__clf), Gradient Boosting (lgbm__\__clf), Random Forest (rf__\__clf) and Neural networks (mlp__\__clf). Boostclean is a joint cleaning and classification method, reported as a separate model type.
(a) F1 (categorical), RMSE (numerical)
(b) KL divergence for all columns
Figure 7. Imputation quality on all columns with mixed missingness for diabetes
(a) F1 disparity (categorical), RMSE disparity (numerical)
(b) KL divergence disparity for all columns
Figure 8. Imputation fairness on all columns with mixed missingness for diabetes
@@ 5.1. Correctness of the Predictive Model
Recall from Section 4.1 that miss-forest was the best performing MVI method for diabetes, under all single-mechanism types, and it was the only method that performed comparably to performance of the model trained on clean data under MNAR. According to Figure 6 (a), miss-forest remains one of the best-performing MVM techniques in terms of F1, for all model types. Further, under single-mechanism missingness, auto-ml (DL-based) was only competitive under MCAR and MAR, however, under multi-mechanism missingness, its F1 is comparable to that of miss-forest, for all model types.
Notably, while boostclean was not the best performing under any single-mechanism missingness, it is highly competitive under multi-mechanism missingness, and even outperforms miss-forest and auto-ml for Decision Trees (DT), Logistic Regression (LR) and Gradient Boosting (LGBM), indicating the competitive edge of joint cleaning and training over impute-then-classify, under complex missingness scenarios (Feng, 2023; Le Morvan et al., 2021). Finally, as for single-mechanism missingness, deletion continues to show the worst performance, with the largest effect of nearly a 15% drop in F1 compared to the model trained on clean data for LGBM.
@@ 5.2. Fairness of the Predictive Model
Recall from Section 4.2 that all MVI methods worsened fairness for the RF model for diabetes under single-mechanism missingness, with clustering being the worst-performing for all missingness types. From Figure 6 (b) we see that the effect of MVI on fairness is highly ML model-specific: for example, clustering improves fairness for LR but worsens it for all other model types, including RF. Note that LR has the worst unfairness (highest TPRD) when the model is trained on clean data, even as compared to deletion with other ML model types. miss-forest, datawig, and auto-ml preserved fairness under single-mechanism missingness, and continue to show good performance under multi-mechanism missingness, for all model types.
@@ 5.3. Stability of the Predictive Model
Recall from Section 4.3 that deletion, clustering and boostclean worsened stability under all three single-mechanism missingness types. In Figure 6 (c), deletion and clustering continue to show poor performance for all model types. Interestingly, boostclean (multiple imputation with joint cleaning and training) is competitive under multi-mechanism missingness: it shows higher label stability than all other MVI approaches for the DT, and is competitive with the best performing imputers for MLP and LGBM.
Simple imputers like median-mode and median-dummy— which were competitive on single-mechanism missingness — are no longer competitive under multi-mechanism missingness, while ML- and DL-based approaches like miss-forest, auto-ml and datawig continue to show good stability.
5.4. Imputation Quality and Fairness
Since multi-mechanism missingness has never been empirically evaluated before (to the best of our knowledge), we additionally report performance on imputation quality metrics in Figure 7. We do not include boostclean since it does not provide imputed values, which we could use to assess imputation quality. In line with  Shadbahr et al. (2023), we find that distributional metrics are more indicative of downstream performance than discrepancy metrics.
deletion, median-mode, and clustering all perform poorly on both discrepancy and distributional metrics, and are also the worst performing in terms of downstream model correctness (for F1). datawig, however, has high error on F1 and RMSE, but not on KL, and this is indicative of its good performance of the downstream predictive model. Further, as part of this work we assert that correctness of downstream models should be evaluated alongside their fairness and stability. We find that imputation quality metrics — especially discrepancy metrics — are not correlated with fairness and stability downstream.
On imputation fairness, auto-ml and miss-forest show the lowest disparity, but they show residual predictive unfairness, as shown for Figure 6. Further, miss-forest has better imputation fairness than median-mode, but leads to worse unfairness of the downstream models. For example, as can be seen in Figure 8 (b), miss-forest has much better imputation fairness than median-mode (0.20 F1 score difference). Yet, as demonstrated in Figure 6 (b), median-mode leads to better fairness of the RF model compared to miss-forest (0.03 TRPD). These observations, once again, corroborate what we observed for for single-mechanism missingness (Section 4.4.3): assessing imputation fairness without considering the fairness of the downstream model is often insufficient.
6. Missingness Shift
Next, we evaluate the effect of missingness shift on the correctness, fairness, and stability of the downstream model. For that, we vary missingness mechanisms and missingness rates in train and test sets separately, first by holding the fraction of nulls in the test set constant (at 30%) and varying the fraction of nulls in the train set (10%, 30% and 50%). Then, we hold the fraction of nulls in the train set constant (at 30%) and vary the fraction of nulls in the test set ($10\%$, $20\%$, $30\%$, $40\%$ and $50\%$). Note that we have fewer settings for train set fractions because varying the test set is less computationally demanding (see Section 3.4).
We report results on diabetes. See Appendix D and GitHub repository for additional results on other datasets, with fixed and variable train and test missingness fractions.
6.1. Correctness of the Predictive Model
@@
Figure 9. F1 of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Figure 10. F1 of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates performance of the model trained on clean data.
Train set missingness. As expected, according to Figure 9 the F1 of the predictive model depends both on the fraction of train nulls, and the missingness mechanism. Under MCAR, for low rates of missingness (10% and 30%), the F1 is constant, and then starts decreasing with more nulls (50%). Under MNAR, with deletion the decrease in F1 is monotonic, and this is in line with conventional wisdom. Under MNAR, the missingness depends on the value itself. This means that when values are missing, they cannot be inferred by other values in the observed data. For this reason, training under MNAR is very sensitive to the train missingness rate.
Interestingly, correctness of ML and DL methods (miss-forest, auto-ml, datawig), and of multiple imputation (boostclean) is relatively insensitive to train set missingness rate. Even at 50% error these methods have sufficient data to learn the missingness pattern reasonably well. This is a striking result given that diabetes was the smallest dataset in our study with only 1,000 samples. Another interesting observation is that clustering is the only method that improves on F1 as the fraction of nulls in the train set increases, for all types of missingness as well as under missingness shift. However, at high rates of training error, even simple imputers like median-mode and median-dummy outperform clustering.
Test set missingness. Figure 10 shows the F1 of the RF model on diabetes, as a function of test set error rate under different single-mechanism missingness (S1-3) and missingness shift (S4-9). We observe that most methods are relatively robust to change in test missingness rates, including also under missingness shift. An important exceptions are scenarios S5 and S7 (Table 1), when train data is missing under MCAR or MAR, while test data is missing under MNAR. For these scenarios, most MVI methods show a steady decrease in F1 with increasing test set missingness rate.
miss-forest is very robust, and shows comparable F1 for all missingness scenarios (including shifts) and test missingness rates. On the other hand, deletion is the most sensitive to the test error rate, and shows the poorest performance on MNAR, even without missingness shift. Our results validate conventional wisdom that MCAR and MAR are easier to model and impute than MNAR and that, in general, increasing the rate of missingness in the test set decreases predictive performance (Shadbahr et al., 2023).
6.2. Fairness of Predictive Model
Figure 11. True Positive Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates performance of the model trained on clean data.
Train set missingness. Figure 11 shows that all methods are sensitive to the train missingness rate when it comes to fairness. For example, miss-forest has stable performance under MNAR train. However, for MCAR and MAR training, the most fair setting is when the error rate in train and test is the same (30%). For larger (50%) and smaller (10%) train error rates, the model shows increased disparity (unfairness). Most ML and DL-based approaches such as auto-ml, datawig, and miss-forest exhibit this behavior, indicating that parity in train and test error rates is one of the most important factors for fairness.
Test set missingness. Figure 12 demonstrates that the trend depends most strongly on test missingness type: deletion under MAR improves fairness as the test error rate increases (for the same reason discussed in Section 4), whereas under MNAR it steadily worsens it. Under MCAR, fairness decreases when the test error rate exceeds the train error rate (30% in train and 50% in test). The miss-forest method is highly robust, with fairness worsening only marginally (by 2%) under MNAR test missingess. In contrast, ML and DL-based methods like clustering, datawig and auto-ml show worse fairness with an increase in the test error rate.
6.3. Stability of Predictive Model
Train set missingness. Figure 30 in Appendix D.1 illustrates that, under MCAR and MAR train missingness, most methods with the exception of boostclean, deletion, and clustering show good stability, and are relatively insensitive to the increase in the train error rate. deletion shows the largest drop in stability with an increase in the train error rate, and this is in line with conventional wisdom as dropping rows with nulls decreases the effective train set size, and less training data makes models more unstable.
Techniques like datawig and boostclean show worse stability under high train error rates, and this effect is the most pronounced under MNAR. miss-forest is, once again, highly robust, and maintains good stability (comparable to that of the model trained on clean data) under all missingness scenarios, rates and shifts.
@@ 6.3. Stability of Predictive Model
Test set missingness. Figure 31 in Appendix D.1 shows that the test missingness rate has very little effect on the stability, for most MVI approaches, with the exception of clustering, which shows higher instability at higher rates of test error, with the largest effect in scenario S7 (MAR train, MNAR test set).
@@ 7. Summary of Experimental Findings
Deletion is the worst strategy. Building on evidence in the scientific community(Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our findings indicate that, in addition to hurting predictive performance, deletion also worsens stability and fairness, arguing against the use of this method in practice.
@@
There is conflicting evidence in the literature on the performance of multiple imputation (MI), and our empirical findings are also mixed and somewhat unexpected. Feng (Feng, 2023) and Le Morvan et al. (Le Morvan et al., 2021) argue that MI has superior predictive performance compared to impute-then-classify.  Graham (2009) find that MI performs well even with less data and small error rates. In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets, and is highly unstable. This can likely be explained by the complexity-stability trade-offs: MI uses a more complex model class and is able to minimize the empirical loss, but due to this very complexity it shows a large variation in predictions under small train set perturbations.
Best performing methods on single- and multi-mechanism missingness are different.
A key novelty in our study is the evaluation under multi-mechanism missingness. We find that MVI methods such as boostclean, which was not competitive for single-mechanism missingness, performs well under mixed or multi-mechanism missingness — which is a more complex missingness pattern and a more realistic scenario likely to be encountered in practice.
Fairness is highly mechanism-specific.
Zhang and Long (2021b) find that different missing mechanisms lead to different prediction fairness when the imputation method is fixed, meaning that an MVI that leads to model fairness under, say, MAR may not lead to fairness under MNAR. We also observe this for all the datasets in our study.
@@
We find that, for large datasets, even simple statistical imputers can preserve stability. In contrast, for small datasets, only a few MVI techniques preserve stability, while deletion, statistical imputation and more complex ML-based approaches all worsen stability.
Sensitivity to missingness rates.
For predictive performance, models are more sensitive to missingness during test than during train. Fairness is highly sensitive to both train and test missingness rates, and, for most methods, the model is most fair when there is the same rate of missingness in both. There is little effect of increasing test missingness on stability, but most imputers show worse stability at higher train missingness rates.
@@
MNAR is theoretically the hardest setting to model, and we empirically find that most state-of-the-art MVI methods do not attain predictive performance of the model trained on clean data. Further, our analysis of missingness shift indicates that training under MNAR is highly sensitive to missingness rates, and even methods that are trained on MNAR but evaluated on “simpler” test missingness (MCAR and MAR) perform poorly. Notably, under multi-mechanism missingness, seeing a small amount of MCAR and MAR data during training boosts MVI performance. Designing MVI techniques that are effective under MNAR is an open research direction.
Imputation quality and fairness metrics are often insufficiently predictive of the correctness, fairness and stability of downstream models.
Shadbahr et al. (2023) showed that distributional imputation quality metrics are a more reliable indicator of the model’s predictive performance, than discrepancy metrics under single-mechanism missingness. Our findings validate this result for multi-mechanism missingness as well. Further, we uncover trade-offs between predictive performance, fairness and stability. Finally, we find that imputation fairness metrics are not predictive of model fairness.
@@
Prior experimental work (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) concluded that there is no definitive answer to the question of which missing value imputation technique is best, according to predictive performance. Our findings support this, and further demonstrate trade-offs between model fairness and accuracy, and model stability and accuracy. This is not bad news, but rather an acknowledgement of the inherent complexity of the task of learning from incomplete data, and a call for more rigorous and holistic evaluation protocols— of the kind demonstrated in this work — in order to select the most suitable MVI technique for a given task.
8. Conclusions and Future Work
Conclusions. We presented Shades-of-Null, a benchmark for responsible missing value imputation. The key novelty of our work is in evaluating predictive performance of models trained on imputed data alongside fairness and stability. Another novelty is our modeling of realistic missingness scenarios that go beyond Rubin’s classic settings of MCAR, MAR and MNAR to include multi-mechanism missingness and missingness shift. We showcased results of an extensive study, with 20,952 experimental pipelines, that uses our benchmark to evaluate performance of a wide range of missing value imputation methods in realistic scenarios. We observed interesting trends, and identified trade-offs between evaluation metrics.

### v2

@@ 4.4. Imputation Quality and Fairness
In Figure 6, we report the imputation quality of different MVI techniques according to F1 score (for categorical columns), RMSE (for numerical columns), and KL divergence (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the F1 of the downstream model.
Imputation Quality.
We find that MVI techniques with vastly different imputation quality can give rise to models with comparable F1 scores, indicating that imputation correctness is not sufficiently predictive of model correctness (Zhang and Long, 2021a). For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; clustering, datawig, median-mode and deletion have an F1 score (imputation) between 0.5-0.6, and auto-ml and miss-forest have an F1 score (imputation) close to 1, but the models trained with all of these techniques have an F1 close to 1. This trend is consistent across all datasets and missingness types, and for numerical columns as well (see Figure 6(b)).
We observe similar trends for KL divergence (see Figure 6(c)), indicating that neither discrepancy nor distributional imputation quality metrics are sufficiently indicative of downstream model performance, refuting Shadbahr et al. (2023)’s claim.
(a) F1 score difference (imputation) vs TPRD (model)
(b) RMSE difference (imputation) vs TPRD (model)
(c) KL divergence difference (imputation) vs TPRD (model)
Figure 7. Imputation fairness vs model fairness: imputation fairness (F1 difference, RMSE difference and KL divergence difference) may not be indicative of model fairness (TPRD).
Imputation fairness.
In Figure 7, we report the imputation fairness of different MVI techniques, according to F1 score difference (for categorical columns), RMSE difference (for numerical columns), and KL divergence difference (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the fairness of the downstream model, according to TPRD. We find that, while model fairness is generally agnostic to missingness type (as discussed in Section 4.2), imputation fairness is highly sensitive to missingness mechanism. For example, in Figure 7(c), miss-forest has good imputation fairness on german under MAR (KL difference of -0.4) but significant imputation unfairness under MCAR (KL difference of 2.25), MNAR (KL difference of 1.1) and mixed missingness (KL difference of 1.4).
Further, imputation fairness is insufficiently predictive of model fairness. For example, on german under mixed missingness, median-dummy has KL difference close to 0, datawig and miss-forest have KL difference between 1 and 1.5, and median-mode and clustering have KL difference between -1 and -1.5, but the models trained using all five of these techniques have a TPRD close to -0.02. Conversely, on diabetes under MAR, auto-ml, miss-forest, median-dummy and clustering all have near perfect imputation fairness (KL difference close to 0), but different model fairness (TPRD between 0.04 and 0.12). We see similar trends for other imputation fairness metrics such as F1 score difference (Figure 7(a)) and RMSE difference (Figure 7(b)).
5. Missingness Shift
Next, we evaluate the correctness, fairness, and stability of MVI techniques under missingness shift. We simulate missingness shift in two ways: (i) by varying the missingness mechanism between training and test (S4-9 in Table 1); and (ii) by varying the missingness rates between training and test. First, we hold the fraction of nulls in the test set constant (at 30%) and vary the fraction of nulls in the training set (10%, 30% and 50%). Then, we hold the fraction of nulls in the training set constant (at 30%) and vary the fraction of nulls in the test set ($10\%$, $20\%$, $30\%$, $40\%$ and $50\%$). Note that we have fewer settings for training missingness rates because varying the test set is less computationally demanding (as discussed in Section 3.4). We discuss results on diabetes, and defer results on other datasets, with fixed and variable training and test missingness rates, to Appendix D.
@@ 5.1. Correctness of the Predictive Model
Figure 8. F1 of the Random Forest model on diabetes, as a function of training set missingness rate. Dashed line shows performance of the model trained on clean data.
Training set missingness. Figure 8 shows the F1 of the Random Forest model on diabetes as a function of training missingness rate. Of all MVI techniques, deletion is most strongly affected by missingness: F1 degrades with increasing missingness rate, and this effect is strongest under MNAR. This includes when MNAR is encountered both during training (the bottom row in Figure 8 shows the steepest decline in F1 compared to other rows—training missingness) and during testing (the right-most column in Figure 8 has the lowest F1 compared to other columns —test missingness).
All other MVI techniques, including boostclean, are generally robust to higher training missingness rates, and only show a 5% decrease in F1 (compared to 10% decrease with deletion), even at rates as high as 50%. A notable exception is clustering, which, surprisingly and somewhat counter-intuitively, has higher F1 at higher training missingness rates, and is actually better under MNAR than under MAR and MCAR, for all missingness rates and scenarios.
Test set missingness.
Figure 45 in Appendix D.2 shows the F1 of the Random Forest model on diabetes as a function of test missingness rate. We find that F1 generally decreases with an increase in test missingness, corroborating the findings of  Shadbahr et al. (2023). A notable exception is miss-forest, which is robust to both forms of missingness shift (i.e., changing missingness mechanisms and missingness rates). As for training set missingness, F1 decreases with an increase in test missingness most steeply under MNAR (both during training and test). And, once again, clustering is an exception to this trend, instead showing invariance to test missingness rates under MNAR train (irrespective of test missingness) but not under MCAR and MAR train.
@@ 5.2. Fairness of the Predictive Model
Training set missingness. Figure 9 shows TPRD of Random Forest on diabetes as a function of training set missingness rate. While we previously found that fairness is generally agnostic to missingness type when it is the same between training and test sets (see Section 4.2), we find that model fairness is highly sensitive to missingness shift — in terms of both different missingness rates and different missingness mechanisms between training and test.
Worryingly, there is no consistent trend across MVI technique, missingness type and training test missingness rate. For example, consider miss-forest, which was the most robust to missingness shift according to F1. Under MCAR training, miss-forest preserves fairness of the model trained on clean data (shown with a dashed grey line) when training and test missingness rates match (at 30% training error rate, fixed at 30% for this experiment), but worsens fairness (higher TPRD) when they are different (at 10% and 50% training error rates). Under MAR training, however, we see the opposite behavior, with miss-forest preserving clean model fairness at 10% and 50% train missingness rates, but worsening fairness when training and test missingness rates are equal (at 30%). Under MNAR training, TPRD increases with an increase in training missingness rate under MCAR and MAR test, and remains constant when there is no shift in missingness mechanism (under MNAR test).
Figure 9. True Positive Rate Difference of Random Forest on diabetes, as a function of training missingness rate. Dashed line shows performance of the model trained on clean data.
Test set missingness.
We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to shifts in test missingness. Figure 46 in Appendix 45 shows the TPRD of the Random Forest model on diabetes as a function of test missingness rate. As before, fairness is highly sensitive to shifts in test missingness. For boostclean, datawig and auto-ml, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic. miss-forest is particularly robust, and although it only matches the fairness of the model trained on clean data under MCAR training, it is agnostic to increases in test missingness rate, under all missingness scenarios. For simpler methods such as deletion, clustering, median-mode and median-dummy, there is no consistent trend, even when missingness types remain the same and only missingness rates change between train and test. For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
In summary, our results corroborate the findings of Guha et al. (2024), and are a cause for concern as they indicate that the MVI techniques that perform best during development may not preserve fairness post-deployment, where shifts are likely to occur.
@@ 5.3. Stability of the Predictive Model
Figure 10. Label Stability of Random Forest on diabetes, as a function of training set missingness rate. Dashed line shows performance of the model trained on clean data.
Training set missingness. Figure 10 shows Label Stability of the Random Forest model on diabetes as a function of training set missingness rate. Under MCAR and MAR missingness, most MVI techniques with the exception of boostclean, deletion, and clustering show good stability (comparable to the model trained on clean data), and are generally insensitive to missingness rates. deletion and clustering are the least stable methods and show a monotonic decrease in stability with increase in missingness rate, with strongest effect under MNAR. auto-ml and miss-forest are the most robust MVI techniques and preserve stability of the clean model under all missingness settings and error rates.
Test set missingness.
@@
Figure 47 in Appendix 46 shows the Label Stability of the Random Forest model on diabetes as a function of test missingness rate. We find that missingness rate has little effect on model stability under all missingness settings, with the exception of clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
@@ 6. Summary of Experimental Findings
Do not drop your nulls! Building on evidence in the scientific community (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our findings indicate that, in addition to hurting predictive performance, deletion also worsens stability and fairness, arguing against the use of this method in practice.
@@
There is conflicting evidence in the literature on the performance of multiple imputation (MI) (Joost R. van Ginkel and van der Voort, 2020), and our empirical findings are also mixed and somewhat unexpected.  Feng (2023) and  Le Morvan et al. (2021) argue that MI has superior predictive performance compared to impute-then-classify approaches.  Graham (2009) and McNeish (2017) find that MI performs well even with less data and small error rates. In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets, and is unstable compared to simpler MVI techniques. This can likely be explained by the complexity-stability trade-offs (Domingos, 2000): MI uses a more complex model class and is able to minimize the empirical loss, but due to this very complexity it shows a large variation in predictions under small train set perturbations.
Fairness is highly missingness-specific.
We find that no MVI technique is consistently fairness-preserving, corroborating the findings of  Zhang and Long (2021b) and Guha et al. (2024). Further, we find that fairness is highly sensitive to changes in missingness rates and missingness mechanisms between training and test sets, which are likely to occur in practice, and therefore a cause for ethical concern.
Imputation quality and fairness metrics are often insufficiently predictive of the correctness and fairness of downstream models.
Shadbahr et al. (2023) posit that distributional imputation quality metrics are a more reliable indicator of a model’s predictive performance than discrepancy metrics. We find that neither discrepancy nor distributional metrics are sufficiently indicative of downstream model performance (according to F1), indicating that poor imputation quality can be overcome by a strong learner. Further, we find that imputation fairness metrics are not predictive of model fairness: perfect imputation fairness can result in downstream model unfairness. Conversely, models trained using MVI techniques with imputation unfairness can still have good downstream fairness.
(a) MCAR missingness
(b) MAR missingness
(c) MNAR missingness
(d) Mixed missingness
Figure 11. Spearman correlation ($\rho$) between MVI technique, model type, and performance metrics (F1, fairness and stability) for different train missingness mechanisms (subplots). TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$. Supplementary plots for each combination of training and test missingness are in Appendix E.
@@
We find that for large datasets even simple statistical imputers can preserve stability. In contrast, for small datasets, only a few MVI techniques preserve stability, while deletion, statistical imputation and more complex ML-based approaches all worsen stability.
Sensitivity to train and test missingness rates.
The predictive performance of models (according to F1) is more sensitive to missingness rates during testing than during training. Fairness, on the other hand, is highly sensitive to both training and test missingness rates. Lastly, there is little effect of test missingness rates on model stability, but most MVI techniques are more unstable at higher training missingness rates.
@@
MNAR is theoretically the hardest setting to model. Empirically, we find that MVI techniques perform poorly with MNAR during testing. Conversely, even methods that are trained on MNAR but evaluated on “simpler” test missingness (MCAR and MAR) perform poorly. Further, we find that model F1 and stability are more sensitive to missingness rates under MNAR than under MCAR and MAR. Designing MVI techniques that are effective under MNAR is an open research direction.
Deep learning outperforms tree-based models on large datasets.
There has been an explosion of interest in deep learning for tabular data (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022). Recall that in Figures 3, 4, and 5 we reported results only for the best-performing models (in terms of F1), to streamline exposition. We find that tree-based methods like random forest and gradient boosted trees still outperform new deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (such as diabetes and german), corroborating the findings of Grinsztajn et al. (2022) and Shwartz-Ziv and Armon (2022), on tabular datasets with (simulated) missingness. GANDALF, however, is the best-performing model on larger datasets (heart and folk-employment in our study), underscoring the growing effectiveness of deep learning on a variety of data modalities.
Best performing approaches under single- and multi-mechanism missingness can be different.
A key novelty in our study is the evaluation under mixed or multi-mechanism missingness— which is a more complex but realistic missingness pattern, likely to be encountered in practice. We find that trends are different under single- and multi-mechanism missingness. Figure 11 shows that evaluation metrics and model types such as gandalf__\__clf and lgbm__\__clf are uncorrelated under single-mechanism missingness, but correlated under mixed missingness, for example for F1 ($\rho$=-0.24 for gandalf__\__clf and $\rho$=-0.37 for lgbm__\__clf) and fairness according to TPRD ($\rho$=0.55 for gandalf__\__clf and $\rho$=-0.4 for lgbm__\__clf).
@@
Prior experimental work (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) concluded that there is no definitive answer to the question of which missing value imputation technique is best according to predictive performance. Our findings support this and demonstrate trade-offs between F1, fairness and stability, based on the architecture of the predictive model. For example, from Figure 11, under multi-mechanism missingness, logistic regression (lr__\__clf) is positively correlated with F1 ($\rho$=0.23) and stability ($\rho$=0.38) but negatively correlated with fairness according to TPRD ($\rho$=-0.23). Random forest (rf__\__clf), on the other hand, is weakly positively correlated with fairness according to TPRD ($\rho$=0.16) and F1 ($\rho$=0.25), but negatively correlated with stability ($\rho$=-0.24). Lastly, deep table-learning (gandalf__\__clf) is strongly positively correlated with fairness according to TPRD ($\rho$=0.55), weakly correlated with stability ($\rho$=0.12), and negatively correlated with F1 ($\rho$=-0.24). This is not bad news, but rather an acknowledgement of the inherent complexity of the task of learning from incomplete data, and a call for more rigorous and holistic evaluation protocols— of the kind demonstrated in this work — to select the most suitable missing value imputation technique and model architecture for a given task.
7. Conclusions, Limitations and Future Work
Conclusions. We presented Shades-of-Null, an evaluation suite for responsible missing value imputation. A key novelty of our work is in evaluating the predictive performance of models trained on imputed data alongside fairness and stability. Another novelty is our modeling of realistic missingness scenarios that go beyond Rubin’s classic settings of MCAR, MAR and MNAR to include multi-mechanism missingness and missingness shift. We showcased results of an extensive study, with 23,940 experimental pipelines, that uses our framework to evaluate a variety of MVI methods in realistic missingness scenarios. We synthesized interesting trends, and identified trade-offs between evaluation metrics.